### v2

@@ Results
Inference time: Once deployed, the trained model relies only on a single network (i.e., the generator $G$ built by a 9-layer ResNet [47]) for reconstruction, making it the fastest among existing hologram reconstruction algorithms. Previous studies [17, 30] have suggested that DL can expedite the reconstruction process through an end-to-end approach. Unlike many DL-based reconstruction algorithms that rely on the multi-height phase recovery (MHPR) approach, LensGAN distinguishes itself by eliminating the necessity for multiple ($I$ number) hologram inputs for reconstruction. Notably, efficient model compression techniques for GANs [48] could further reduce the computational cost by 46×46$\times$46 ×, but we maintain the original model size to avoid resolution degradation. We evaluated LensGAN’s performance on our edge device (see MethodsMethods) using hologram images with $512\times 512$ pixels ($px$). LensGAN was compared against the self-supervised model GedankenNet [17], the supervised model U-Net [37], and the iterative algorithm MHPR [38], as outlined in Table 1. LensGAN achieved an equivalent inference time of $0.0031\,{s\cdot mm}^{-2}$ for a $1\,mm^{2}$ area of our microfluidic chip, showing its exceptional speed and efficiency in hologram reconstruction.
Table 1: Holographic image inference time (for $I$ input hologram(s), $512\times 512\,px$)
@@
00footnotetext: M: million. Evaluations were performed on a single PC with an i9-13900K CPU and an RTX 3090 GPU.
Large FOV imaging of microfluidics
Figure 4: Reconstruction of a microfluidic chip ($30\times 5\,mm$). a, ROI in the hologram, i.e., the functional region containing dynamic droplets. b, Synthesized reconstruction result of the ROI (also see Supplementary video 1). We also analyzed the cross-sectional profiles at the center of 3 patches, labeled $i$, $o$, and $\hat{o}$, as denoted by the white dash line at $i$.
Microfluidic platforms have transformed biomedical research by enabling high-throughput experiments with biological samples down to the scale of single cells or molecules [49]. Despite their increasing adoption, a critical challenge remains. Existing high-resolution imaging techniques cannot simultaneously monitor multiple ROIs on a single chip. For instance, imaging the formation of droplets in droplet-based microfluidics often impedes simultaneous observation of downstream operations, such as droplet sorting or merging, occurring in a different ROI [49].
Our GenLFI approach overcomes this limitation. We demonstrate its capabilities in a microfluidic experiment, generating and dynamically imaging droplets across various chip sections concurrently. Data collection for the training dataset of this microfluidic experiment (see Methods) took approximately 20 minutes. Unlike most LFI systems, this procedure is automated, requiring no manual annotation. We acquired two-domain training data using our LFI system and a conventional bright-field microscope, labeled as $i$ and $o$, respectively (see Fig.1b). These two-domain images, i.e., $i$ and $o$, constituted the ‘GenLFI-Microfluidics’ dataset for LensGAN training and testing. Reconstruction results (see Fig.4) demonstrate that the trained LensGAN can generate high-resolution ($1792\times 768\,px$) bright-field images ($\hat{o}$) across the entire ROI (i.e., spanning a full microfluidic chip).
The results show that our approach not only preserves the global structure from the original hologram to the generated image but also ensures consistent transverse x-coordinate positioning of features, as evident in the re-focusing process shown in Fig.4b. Furthermore, the synthesized large FOV image (Fig.4b) reveals that although the trained LensGAN model infers each patch independently (see Methods), the microfluidic channels are aligned consistently across patches. This suggests the robustness of our approach in handling large-scale reconstruction tasks. This capability can facilitate monitoring and visual-based control for intricate microfluidic systems.
Large FOV imaging of spheroids
@@
Figure 5: Reconstruction of cell spheroids. a, MeT-5A NF2KO spheroid images acquired using different imaging systems: bright-field microscopy with a 10x objective lens (left), digital holographic microscopy (DHM) with a 10x objective lens (center), and our GenLFI platform (right). b, Full-FOV reconstruction by LensGAN on an unseen well, post-processed by global panorama stitching algorithm (see Supplementary Information 6) and local image-stitching [50]. The structures are readily resolvable in the penetrated light parts at the edges of spheroids.
@@
Cell spheroids, as straightforward yet extensively utilized multi-cellular 3D models, emerge from the natural inclination of most adherent cells to cluster together [51]. The imaging of spheroids poses significant challenges due to their considerable thickness, which can reach up to hundreds of microns, and their high density. As noted previously, conventional holographic imaging systems, including LFI systems, involve complex optical fields, resulting in contaminated holograms with overlapping scattered patterns and shadows (see Fig.1a). When utilizing digital holographic microscopy (DHM) with a 10x objective lens and a $635\,nm$ wavelength laser light source, only limited structural and intensity information can be retrieved (see Fig.5a).
With the addition of a light contrast-enhancing polarization filter, a bright-field microscope equipped with a 10x objective lens can discern a partial sectional structure at the periphery by adjusting the focus of the objective lens (Fig.5a). A comparable light penetration capability can also be achieved using our LFI platform by employing a red laser light source with a wavelength of $635\,nm$. Therefore, we integrated inverted bright-field microscopy to acquire data guidance images $o$, facilitating holographic reconstruction of similar quality but across an ultra-large FOV.
@@
Fig.5b shows that our approach can reconstruct the spheroid edge structure and shapes from the measured holograms in a large FOV of an engineered cancer model. Notably, spheroids exhibit lateral structures that might replicate some of the more complex cellular environments, thereby mimicking more relevant scenarios while using high throughput platforms. Leveraging the advantages of the large FOV, our approach proves effective for monitoring the multiparametric spheroid growth patterns, which will likely allow for insights into biomechanical features, cell-cell interactions, and readily quantifiable multiparametric features of spheroid outgrowth assays in a scalable and cost-effective manner.
@@ Discussion
GenLFI, a generative deep learning framework for LFI, eliminates the need for traditional forward imaging models by introducing data-driven guidance through unsupervised learning. This framework enables imaging in dynamic, noisy optical fields, previously not achievable with LFI. GenLFI achieves an over 20 times larger FOV than current LFI systems, enabling real-time monitoring for in vivo (e.g., zebrafish, see Fig.S1 in Supplementary Information) or in vitro (e.g., cell-in-droplets, 3D cell models) studies. Images are reconstructed using a DL network, significantly faster than traditional iterative techniques [22, 23, 52, 53, 38], while achieving sub-pixel resolution without requiring adjustments to the light source. These advancements demonstrate the potential of leveraging generative AI to overcome optical limitations.
The dynamic nature of our imaging scenarios makes it impractical to obtain paired ground truth and raw holograms for training. Therefore, GenLFI leverages unsupervised learning algorithms, inspired by those designed for image-to-image (I2I) transformation tasks in computer vision. This approach prioritizes hardware-agnostic and reproducible workflows with simple setups. While the current implementation utilizes a telescope cold camera with a resolution of $5.9\,\mu m/px$, GenFLI is compatible with higher-resolution sensors and multi-angle illumination for potential 3D LFI capabilities.
@@
Our investigation into the physical correlation between holograms and bright-field images achieved leading results based on FID and SSIM metrics. However, challenges remain with such I2I-based approaches, primarily concerning data quality, quantity, and diversity. These factors directly impact reconstruction results, as demonstrated by artifacts observed in reconstructions from limited datasets (see the results from single-sample, paired datasets Fig.2b and Fig.S7 in Supplementary Information). Additionally, training GANs can be challenging due to potential issues like model collapse and non-convergence. While exploring diffusion models [54, 55, 56, 57] as an alternative shows promise, their limitations in handling unpaired and unaligned image translation tasks require further investigation. Furthermore, existing evaluation metrics like FID and KID primarily assess image-level distributions and lack the ability to evaluate the alignments between objects across different image patches. Developing new metrics accurately reflecting alignment between different images captured from the same objects is crucial for future advancements.
The combination of GenLFI and large vision models (LVM) [58, 59] holds promise for achieving a more generalized reconstruction capability for various samples and light field conditions. In addition, leveraging processors with advanced and specified architectures [60, 61, 62] could enable real-time image reconstruction with these large models.
@@
Despite limitations due to optical conditions, GenLFI’s high-throughput capability opens doors for various biomedical applications. GenLFI offers a large FOV, enabling rapid analysis of large biological samples with minimal equipment. This capability is particularly valuable in resource-limited settings lacking traditional histopathology services. By incorporating digital staining through generative deep learning, GenLFI can bring advanced imaging techniques to such settings. Furthermore, GenLFI’s cost-effective and compact setup facilitates routine monitoring of cell culture and spheroid growth directly within incubators, reducing phototoxicity while enabling continuous monitoring of dynamic cellular processes across a large FOV. The testing of drug libraries based on dynamic processes, utilizing single-cell/cell models or droplet-based microfluidic systems, can now be obtained simultaneously from microplate-like platforms, paving the way for developing next-generation, autonomous, high-throughput drug screening systems.
@@ Methods
GenLFI utilizes a versatile optical setup for capturing holograms. The optical setup comprises a light source, a CMOS sensor, and a test bench to stabilize their positions. We employed two types of light sources, including a blue LED light source for enhanced resolution (used for microfluidics, test chart, beads) and a red laser light source for deeper penetration depth (used for spheroids). For LED illumination, the light source is Thorlabs M420L3 (wavelength $420\,nm$) blue LED, with power supply (15V 1A) and T-Cube LED Driver (Thorlabs LEDD1B). For laser illumination, we used an optical fibre-coupled red laser source (Thorlabs S1FC635, wavelength $635\,nm$). The mechanical setup consists of a cage plate (Thorlabs CP33/M), a mounting bracket (ThorLabs 973/579-7227), a $25\,mm$ linear stage (Thorlabs XR25C) with an end-mounted micrometer, a post holder (Thorlabs LPH200), a post (Thorlabs P350/M), and an adjustable platform (Thorlabs C1519). The telescope camera (ZWO ASI 2400MC Pro) features a full-frame Sony IMX410 CMOS sensor, $36\times 24\,mm$ in size, with a $5.94\,\mu m/px$ resolution. Mounted on an experimental bracket (FAMA MV test stand), the camera aligns with the light source and the imaging object along a single axis. The distance from the light source to the object’s bottom plane is $430\,mm$, and from the object’s bottom plane to the CMOS sensor is $12.5\,mm$.
@@
LensGAN, a core component of GenLFI, reconstructs high-resolution images from captured holograms. LensGAN generally considers the measured and reconstructed holograms ($i$ and $\hat{i}$) and the measured and generated bright-field images ($o$ and $\hat{o}$) into two domains, $X$ and $Y$, respectively. It leverages a GAN framework with a cycle-consistent training strategy to learn the complex mapping between holograms (domain $X$) and corresponding bright-field images (domain $Y$). The architecture of LensGAN comprises 6 networks: generators of reconstruction and reverse mappings $G:X\rightarrow Y$ and $F:Y\rightarrow X$, selectors $\beta_{X}$ and $\beta_{Y}$, and discriminators $D_{X}$ and $D_{Y}$. During training, all 6 networks function following the training loop as shown in Fig.1a. In the testing phase, only the trained generator $G$ is used for reconstruction. The outputs of LensGAN are reconstructed bright-field images $\hat{o}=G(i)$ and generated holograms $\hat{i}=F(o)$.
@@
Generators $G$ and $F$ handle the fundamental image transformation task, while the remaining 4 networks contribute to the training process. $G$ and $F$, each composed of 9 residual blocks (ResNet [47]), feature an image encoder for extracting image characteristics and a decoder for generating images. The selector network $\beta$ [63] is responsible for pairing images based on feature alignments and is constructed with an adaptive average pooling layer, 4 convolutional networks, and one fully connected network (FCN) with a Softmax layer. The discriminator $D$ is designed to differentiate between images, emphasizing structural information and intricate details within the complex amplitude. Recent advancements in Transformers [64] have demonstrated the equivalence of multi-layer and one-layer attention mechanisms [65]. By leveraging this finding, we designed $D$ with a Markovian architecture, merging a one-layer attention module with 5 downsampling convolutional layers to augment amplitude details while maintaining minimal computational overhead (see Supplementary Information 4, 5).
Combining these networks, the training loop adheres to the cycle structure [66], guided by several loss functions, depicted in Fig.1b and Fig.S2 in Supplementary Information. To augment phase structural feature information during the training, we devised a phase contrast module positioned to constrain the physical correlation between $X$ and $Y$ domains. The observed samples in LFI are generally light-permeable, resulting in duotone holograms captured under a single-wavelength light source. In such scenarios, phase contrast [6] and dark-field microscopes are commonly employed, where a phase ring and other lenses are combined to create a phase contrast. Inspired by this technique, our approach is to develop a module that learns the phases of the two domains contrastively, thereby mimicking a similar mechanism and enhancing the physical correlations between the object’s hologram and bright-field images. Technically, we create the phase-contrast loss function $\mathcal{L}_{\mathrm{PC}}$ by a multi-scale structural similarity index measure (MS-SSIM) [67] for the amplitude spectra:
$\min_{G}\mathcal{L}_{\mathrm{PC}}(X,Y)=\mathbb{E}_{i\sim P_{X}}\left[1-\operatorname{MS-SSIM}(|\mathcal{FT}(i)|,|\mathcal{FT}(\hat{o})|)\right],$ (3)
@@
where $|\mathcal{FT}(\cdot)|$ denotes the magnitudes of the individual frequency components obtained from the Fourier transform ($\mathcal{FT}$) of an image, and $P_{X}$ is the probability distribution for $i$. $\operatorname{MS-SSIM}$ is defined as:
@@
$\operatorname{MS-SSIM}(x,y)=\left[l_{M}(x,y)\right]^{\gamma_{M}}\cdot\prod_{j=1}^{M}\left[c_{j}(x,y)\right]^{\delta_{j}}\left[s_{j}(x,y)\right]^{\epsilon_{j}},$ (4)
@@
where $x$ and $y$ and two input image patches, $M=5$ denotes the highest scale number set for the system to iteratively apply a low-pass filter between $x$ and $y$. $\mu$ and $\sigma^{2}$ stand for mean value and variance, respectively, which are implemented by 2D convolutions. $l(x,y)$, $c(x,y)$, and $s(x,y)$ represent luminance, contrast, and structure comparison measures between $x$ and $y$, with control weights $\delta_{1}=\epsilon_{1}=0.0448$, $\delta_{2}=\epsilon_{2}=0.2856$, $\delta_{3}=\epsilon_{3}=0.3001$, $\delta_{4}=\epsilon_{4}=0.2363$, and $\gamma_{5}=\delta_{5}=\epsilon_{5}=0.1333$. LensGAN enforces all inputs as 8-bit images with a pixel value range of 0 to 255; the constants are set as $C_{1}=(0.01\times 255)^{2}$, $C_{2}=(0.03\times 255)^{2}$, and $C_{3}=C_{2}/2$.
Note that in (3), we added $\mathcal{L}_{\mathrm{PC}}$ only for the mapping $G:X\rightarrow Y$, the transformation process of holographic reconstruction, with a control weight $\tau_{\mathrm{PC}}$. Similarly, only for $G$, selector $\beta_{X}$ is regularized by the effective sample size loss [63] for $\beta_{X}$:
$\min_{\beta_{X}}\mathcal{L}_{\mathrm{ESS}}(X)=\|\beta_{X}\|_{2},$ (6)
@@
For the whole cycle (reconstruction mapping $G:X\rightarrow Y$ and reverse mapping $F:Y\rightarrow X$), there is a GAN loss function $\mathcal{L}_{\mathrm{GAN}}$ to constrain $G$, $F$ and $D$, with two consistency loss functions $\mathcal{L}_{\mathrm{cyc}}$ and $\mathcal{L}_{\mathrm{idt}}$ to further regularize $G$ and $F$ [68, 66, 63]. They are reweighted least squares adversarial loss
@@
reweighted cycle consistency loss
@@
and reweighted identity loss

### v4

@@ Results
Speed: Previous studies [20, 36] have suggested that DL can expedite the reconstruction process through a pure end-to-end approach. Compared to previous DL approaches, LensGAN’s capability for dynamic imaging is enhanced by eliminating the need for multiple ($I$ number) hologram measurements, sub-pixel super-resolution, and Fourier phase and amplitude pre-processing. Additionally, its fast model reference speed further contributes to its efficiency. Once deployed, the trained model relies only on a single network (i.e., the generator $G$ built by a 9-layer ResNet [68]) for direct reconstruction from the input. We evaluated LensGAN’s performance (see Methods) using $512\times 512$ pixels ($px$) hologram images. LensGAN was compared against the self-supervised model GedankenNet [20], the supervised model U-Net [69], and the iterative algorithm MHPR [32], as outlined in Table 1. LensGAN achieved an equivalent inference time of $0.0031\,{s\cdot mm}^{-2}$ for a $1\,mm^{2}$ area of our microfluidic chip (see Fig.5), showing its exceptional speed and efficiency in hologram reconstruction. Note also that efficient model compression techniques for GANs [70] can further reduce the computational cost by 46×46$\times$46 ×, but we maintain the original model size to avoid resolution degradation.
Table 1: Holographic reconstruction speed (for $I$ input raw hologram(s) in $512^{2}\,px$)
@@
00footnotetext: * M: million. Reconstruction time refers to the total duration from raw hologram input through phase retrieval to model inference. Inference time is measured from the retrieved phase input to the model output, excluding the phase retrieval step (i.e., Fourier transform). All evaluations were conducted on a single PC with an i9-13900K CPU and an RTX 3090 GPU. Note that GedankenNet and MHPR require multiple measurements for input data, and U-Net requires manual annotation for training.
Spheroids monitoring
@@
Figure 4: Reconstruction of unpaired cell spheroids. a, MeT-5A NF2KO spheroid images acquired using different imaging systems: bright-field microscopy with a 10x objective lens (Microscope), digital holographic microscopy (DHM) with a 10x objective lens, our GenLFI platform (LensGAN). With the laser’s high intensity, deeper structural information of the 3D spheroid is penetrated and illustrated. b, Full-FOV reconstruction by LensGAN on an unseen well, post-processed by global panorama stitching algorithm (see Supplementary Information 5) and local image-stitching [71]. The structures are readily resolvable in the penetrated light parts at the edges of spheroids. c, Swelling spheroids were observed over a 2-hour time-lapse, the maximum duration allowable outside the incubator. The growth region is highlighted in the bright areas of the differential image between the two frames.
@@
Cell spheroids, as straightforward yet extensively utilized multi-cellular 3D models, emerge from the natural inclination of most adherent cells to cluster together [72]. The imaging of spheroids poses significant challenges due to their considerable thickness, which can reach up to hundreds of microns, and their high density. As noted previously, conventional holographic imaging systems, including LFI systems, involve complex optical fields (Fig.1a), resulting in contaminated holograms with overlapping scattered patterns and shadows. In this case, both lens-based digital holographic microscopy (DHM) and the Autofocus algorithm [23, 52, 35] cannot reconstruct the 3D feature of spheroids (see Fig.4a).
With the addition of a light contrast-enhancing polarization filter, a bright-field microscope equipped with a 10x objective lens can discern a partial sectional structure at the periphery (Fig.4a, Microscope (10x)). A comparable light penetration capability can also be achieved using our LFI platform by employing a red laser light source with a wavelength of $635\,nm$. Higher laser intensity allows for the extraction of more 3D features from holograms. The results indicate that LensGAN can not only reconstruct 3D structures but also maintain robustness under varying light conditions. This demonstrates a level of resilience to environmental factors that current LFI methods do not achieve.
@@
Fig.4b shows that our approach can reconstruct the spheroid edge structure and shapes from the measured holograms in a large FOV of an engineered cancer model. Notably, spheroids exhibit lateral structures that might replicate some of the more complex cellular environments, thereby mimicking more relevant scenarios while using high throughput platforms. Recent research [73] has shown that time-lapse label-free monitoring is an effective and reliable evaluation tool for evaluating the effects of gene editing on organoids. Following this strategy, we recorded 2-hour time-lapse holograms and reconstructed them using LensGAN. The frame difference highlights the growth distribution on the spheroid, which has not been shown in such a large FOV in real-time before. Leveraging the advantages of the large FOV, our approach proves effective for monitoring the multiparametric spheroid growth patterns, which will likely allow for insights into biomechanical features, cell-cell interactions, and readily quantifiable multiparametric features of spheroid outgrowth assays in a scalable and cost-effective manner.
Microfluidics screening
Figure 5: Reconstruction of a microfluidic chip ($30\times 5\,mm$), scale bar: $200\,\mu m$. a, ROI in the hologram, i.e., the functional region containing dynamic droplets. b, Synthesized reconstruction result of the ROI (also see Supplementary video 1). We also analyzed the cross-sectional profiles at the center of 3 unpaired patches, labeled $i$, $o$, and $\hat{o}$, as denoted by the red dash lines in the middle.
Microfluidic platforms have transformed biomedical research by enabling high-throughput experiments with biological samples down to the scale of single cells or molecules [74]. Despite their increasing adoption, a critical challenge remains. Existing high-resolution imaging techniques cannot simultaneously monitor multiple ROIs on a single chip. For instance, imaging the formation of droplets in droplet-based microfluidics often impedes simultaneous observation of downstream operations, such as droplet sorting or merging, occurring in a different ROI [74]. Most current LFI methods require multiple measurements from different positions and are unable to capture real-time microfluidic droplets. The only viable approach is forward optical modeling. However, this approach is impractical for dynamic objects that change shape.
Our GenLFI approach overcomes these limitations. We demonstrate its capabilities in a microfluidic experiment, generating and dynamically imaging droplets across various chip sections concurrently. Data collection for the training dataset of this microfluidic experiment (see Methods) took approximately 20 minutes. Unlike most LFI systems, this procedure is automated on random ROIs, requiring no manual annotation. We acquired two-domain training data using our LFI system and a conventional bright-field microscope, labeled as $i$ and $o$, respectively (see Fig.1b). These two-domain images, $i$ and $o$, constituted the ‘GenLFI-Microfluidics’ dataset for LensGAN training and testing. Reconstruction results (see Fig.5) demonstrate that the trained LensGAN can generate high-resolution ($1792\times 768\,px$) bright-field images ($\hat{o}$) across the entire ROI, covering a full microfluidic chip.
The results show that our approach not only preserves the global structure from the original hologram to the generated image but also ensures consistent transverse x-coordinate positioning of features, as evident in the re-focusing process shown in Fig.5b. Furthermore, the synthesized large FOV image (Fig.5b) reveals that although the trained LensGAN model infers each patch independently (see Methods), the microfluidic channels are aligned consistently across patches. This suggests the robustness of our approach in handling large-scale reconstruction tasks. This capability can facilitate monitoring and visual-based automation for intricate microfluidic systems.
@@ Discussion
GenLFI, a generative physics-informed deep learning framework for LFI, eliminates the need for traditional forward imaging models by introducing random training data through unsupervised learning and physics constraints and regularization. This framework allows imaging in dynamic, noisy optical fields, which was previously unattainable with model-based LFI using model-based methods. For 3D samples, in a single shot with sub-pixel resolution, GenLFI achieves an over 20 times larger FOV than current LFI systems [75, 36], enabling real-time monitoring for in vivo (e.g., zebrafish, see Fig.S1 in Supplementary Information) or in vitro (e.g., cell-in-droplets, 3D cell models) studies. GenFLI is compatible with higher-resolution sensors and multi-angle illumination for potential 3D LFI capabilities. Images are reconstructed using a DL network, significantly faster than traditional iterative techniques [28, 29, 76, 77, 32], without requiring multi-position measurements and data pre-processing. These advancements demonstrate the potential of leveraging generative AI to overcome optical limitations.
@@
Our investigation into the physical correlation between holograms and bright-field images achieved leading results based on various metrics. However, challenges remain with such I2I-based approaches, primarily concerning data quality, quantity, and diversity. Training GANs can be challenging due to potential issues like model collapse and non-convergence. While exploring diffusion models [78, 79, 80, 81] as an alternative shows promise, their limitations in handling unpaired and unaligned image translation tasks require further investigation.
It is notable that, rather than a low-cost replacement for traditional lens-based optical systems, we consider LFI systems as high-content imaging tools for data-driven biomedical applications. For instance, a modest FDA-approved drug library can contain around 2000 samples and 2 million or more pairwise combinations [82]; GenLFI could utilize the repetitive nature of these samples’ large-scale high-content data. As GANs lack generalization abilities, LensGAN, a very initial exploration under the GenLFI framework, is more suitable for large-scale screening automation on similar modality samples. In the future, the combination of GenLFI and foundation large vision models (LVM) [83, 84] holds promise for achieving a more generalized reconstruction capability for various samples and light field conditions. In addition, leveraging chips with advanced and specified architectures [85, 86, 87] and cloud platforms could enable real-time image reconstruction with these large models.
@@
Despite limitations due to optical conditions, GenLFI’s high-throughput capability opens doors for various biomedical applications. GenLFI offers a large FOV, enabling rapid analysis of large biological samples with minimal equipment. Furthermore, GenLFI’s cost-effective and compact setup facilitates routine monitoring of cell culture and spheroid growth directly within incubators, reducing phytotoxicity while enabling continuous monitoring of dynamic cellular processes across a large FOV. The testing of drug libraries based on dynamic processes, utilizing single-cell/cell models or droplet-based microfluidic systems, can now be monitored in parallel, paving the way for developing next-generation, autonomous, high-throughput drug screening systems.
@@ Methods
GenLFI utilizes a versatile optical setup for capturing holograms. The optical setup comprises a light source, a CMOS sensor, and a test bench to stabilize their positions. We employed two types of light sources, including a blue LED light source for enhanced resolution (used for microfluidics, test chart, beads) and a red laser light source for deeper penetration depth (used for spheroids). For LED illumination, the light source is Thorlabs M420L3 (wavelength $420\,nm$) blue LED, with power supply (15V 1A) and T-Cube LED Driver (Thorlabs LEDD1B). For laser illumination, we used an optical fibre-coupled red laser source (Thorlabs S1FC635, wavelength $635\,nm$). The mechanical setup consists of a cage plate (Thorlabs CP33/M), a mounting bracket (ThorLabs 973/579-7227), a $25\,mm$ linear stage (Thorlabs XR25C) with an end-mounted micrometer, a post holder (Thorlabs LPH200), a post (Thorlabs P350/M), and an adjustable platform (Thorlabs C1519). The telescope camera (ZWO ASI 2400MC Pro) features a full-frame Sony IMX410 CMOS sensor, $36\times 24\,mm$ in size, with a $5.94\,\mu m/px$ resolution. Another camera used for tissue imaging is Thorlabs CS895MU, with $14\times 7\,mm$ size and $3.45\,\mu m/px$ resolution. Mounted on an experimental bracket (FAMA MV test stand), the camera aligns with the light source and the imaging object along a single axis. The distance from the light source to the object’s bottom plane is $430\,mm$, and from the object’s bottom plane to the CMOS sensor is $12.5\,mm$.
@@
LensGAN, a core component of GenLFI, reconstructs high-resolution images from captured holograms. LensGAN generally considers the measured and reconstructed holograms ($i$ and $\hat{i}$) and the measured and generated bright-field images ($o$ and $\hat{o}$) into two domains, $X$ and $Y$, respectively. It leverages a GAN framework with a cycle-consistent training strategy to learn the complex mapping between holograms (domain $X$) and corresponding bright-field images (domain $Y$). The architecture of LensGAN comprises 6 networks: generators of reconstruction and reverse mappings $G:X\rightarrow Y$ and $F:Y\rightarrow X$, selectors $\beta_{X}$ and $\beta_{Y}$, and discriminators $D_{X}$ and $D_{Y}$. During training, all 6 networks function following the training loop as shown in Fig.1d and supplementary Fig. S3. In the testing phase, only the trained generator $G$ is used for reconstruction. The outputs of LensGAN are reconstructed bright-field images $\hat{o}=G(i)$ and generated holograms $\hat{i}=F(o)$.
Neural networks
@@
Generators $G$ and $F$ handle the fundamental image transformation task, while the remaining 4 networks contribute to the training process. $G$ and $F$, each composed of 9 residual blocks (ResNet [68]), feature an image encoder for extracting image characteristics and a decoder for generating images. The selector network $\beta$ [48] is responsible for pairing images based on feature alignments and is constructed with an adaptive average pooling layer, 4 convolutional networks, and one fully connected network (FCN) with a Softmax layer. The discriminator $D$ is designed to differentiate between images, emphasizing structural information and intricate details within the complex amplitude. Recent advancements in Transformers [53] have demonstrated the equivalence of multi-layer and one-layer attention mechanisms [88]. By leveraging this finding, we designed $D$ with a Markovian architecture, merging a one-layer attention module with 5 downsampling convolutional layers to augment amplitude details while maintaining minimal computational overhead (see Supplementary Information 4).
Loss functions
Combining these networks, the training loop adheres to the cycle structure [43], guided by several loss functions.
Phase contrast loss: Instead of directly using the Fourier transform (FT) to pre-process the holograms and decode the phase in MHPR [32, 20], we utilize the phase information as a soft learning bias during the learning process.
To augment phase structural feature information during the training, we devised a phase contrast module positioned to constrain and regularize the physical correlation between the hologram ($X$) and bright field ($Y$) domains. In LFI the observed samples are typically transparent or semi-transparent, resulting in holograms that encode phase variations rather than intensity-based structural features. These holograms are typically captured using coherent, single-wavelength illumination, which lacks intrinsic contrast and leads to duotone or low-contrast interference patterns.
Phase contrast microscopy [1] enhances contrast by converting phase shifts from refractive index variations into detectable intensity differences. This is achieved using a phase ring in the objective’s back focal plane to shift the background light’s phase, producing interference that reveals fine structural details, especially in transparent biological samples.
Inspired by this technique, we design a computational analog that introduces a soft inductive bias into the learning process. Specifically, our model includes a phase-contrast loss $\mathcal{L}_{\mathrm{PC}}$ that encourages the generator to preserve and translate frequency-domain phase structures between the hologram and bright-field domains. By applying a multi-scale structural similarity index measure (MS-SSIM [89]) on the amplitude spectra of the Fourier-transformed images, we promote consistency in the spatial frequency content and thereby encourage the model to learn phase-sensitive transformations:
$\min_{G}\mathcal{L}_{\mathrm{PC}}(X,Y)=\mathbb{E}_{i\sim P_{X}}\left[1-\operatorname{Sim}(|\mathcal{F}(i)|,|\mathcal{F}(\hat{o})|)\right]$ (3)
@@
. Here, $|\mathcal{F}(\cdot)|$ denotes the magnitudes of the individual frequency components obtained from the image’s Fourier transform ($\mathcal{F}$), and $P_{X}$ is the probability distribution for $i$. $\operatorname{Sim}$ is the MS-SSIM function applied to the amplitude spectra, defined as:
@@
$\operatorname{Sim}(x,y)=\left[l_{M}(x,y)\right]^{\gamma_{M}}\cdot\prod_{j=1}^{M}\left[c_{j}(x,y)\right]^{\delta_{j}}\left[s_{j}(x,y)\right]^{\epsilon_{j}},$ (4)
@@
where $x$ and $y$ are two input images, $M=5$ denotes the highest scale number set for the system to iteratively apply a low-pass filter between $x$ and $y$. $\mu$ and $\sigma^{2}$ stand for mean value and variance, respectively, implemented by 2D convolutions. $l(x,y)$, $c(x,y)$, and $s(x,y)$ represent luminance, contrast, and structure comparison measures between $x$ and $y$, with control weights $\delta_{1}=\epsilon_{1}=0.0448$, $\delta_{2}=\epsilon_{2}=0.2856$, $\delta_{3}=\epsilon_{3}=0.3001$, $\delta_{4}=\epsilon_{4}=0.2363$, and $\gamma_{5}=\delta_{5}=\epsilon_{5}=0.1333$. LensGAN enforces all inputs as 8-bit images with a pixel value range of 0 to 255; the constants are set as $C_{1}=(0.01\times 255)^{2}$, $C_{2}=(0.03\times 255)^{2}$, and $C_{3}=C_{2}/2$.
This loss not only guides the model to preserve fine structural details that are typically encoded in phase, but also helps prevent unrealistic geometric distortions across domains (see Fig.2a) by enforcing fidelity in high-frequency phase content (see the FRC evaluation detailed in Supplementary Information 10).
Effective sample size loss: we adopted the sample reweighting strategy for selecting the aligned image pairs [90, 48]. Note that in Eq. (3), we added $\mathcal{L}_{\mathrm{PC}}$ only for the mapping $G:X\rightarrow Y$, the transformation process of holographic reconstruction, with a control weight $\tau_{\mathrm{PC}}$. Similarly, only for $G$, the selector network $\beta_{X}$ and $\beta_{Y}$ is regularized by the effective sample size loss [48]:
$\mathcal{L}_{\mathrm{ESS}}(X,Y)=\|\beta_{X}\|_{2}+\|\beta_{Y}\|_{2},\quad\min_{\beta_{X},\beta_{Y}}\mathcal{L}_{\mathrm{ESS}}(X,Y).$ (6)
@@
Selectors $\beta_{X}$ and $\beta_{Y}$ take in real images (i.e., hologram $i$ and bright field $o$) and produce importance weights $\beta_{X}(i)$ and $\beta_{Y}(o)$ for each sample. These weights are used to scale the unweighted generator loss components (e.g., cycle-consistency, adversarial loss) dynamically.
Semantic consistency losses: For the whole cycle (reconstruction mapping $G:X\rightarrow Y$ and reverse mapping $F:Y\rightarrow X$), there is a GAN loss function $\mathcal{L}_{\mathrm{GAN}}$ to constrain $G$, $F$ and $D$, with two consistency loss functions $\mathcal{L}_{\mathrm{cyc}}$ and $\mathcal{L}_{\mathrm{idt}}$ to further regularize $G$ and $F$ [91, 43, 48]. They are reweighted least squares adversarial loss to examine the generator:
@@
reweighted cycle consistency loss:
@@
and reweighted identity loss to encourage the model more conservative [92]: